OpenStudy (anonymous):

What is the positive difference between two solutions of 3 is equal to absolute value of 1 minus 2q?

OpenStudy (haseeb96):

from the given question yur equation becomes 3= 1-2q now solve it simply 2q=1-3 2q=-2 q=-2/2 q=-1 -1 is the positive difference between it because when u will put -1 in the place of q in the above given equation u will get 3 =3 - multiply with - = + If u ahve any more difficulty then ask i will reply u after praying MAgrib Prayer